A typical human with a mass of 70Kg and produces about 2000 Calories of heat t per day, A)the rate of heat in watts? B) also in Cal/h
Answer #1

m=70kg, Q=2000cal/day =2000*4.184J/cal/(86400s) =2000*4.184/86400=0.09685W

B)Q=2000cal/day=2000cal/24h =83.33cal/h